OpenStudy (thesecret20111):

\[(8\sqrt{18}\times3)\div16\sqrt{27}\] simplify showing full working...

OpenStudy (anonymous):

\[24\sqrt2\] \[\div\] \[16\sqrt3\] then reduce the 24 and 16 \[\sqrt 18\] = 9*2 so pull a 3 out and multiply 8*3... follow the same for \[16\sqrt27\]

OpenStudy (anonymous):

One way to think about this is \[ \frac{3(8\sqrt{18})}{16\sqrt{27}} = \frac{3(8\sqrt{18})}{16\sqrt{9*3}} = \frac{3(8\sqrt{18})}{16*3\sqrt{3}} = \frac{8\sqrt{18}}{16\sqrt{3}} = \frac{\sqrt{18}}{2\sqrt{3}} = \sqrt{\frac{18}{12}} = \sqrt{\frac{3}{2}}\]

OpenStudy (thesecret20111):

Thanks guys! :D